Find V1 and V2

4 Ω . 2Α 4Ω s * V2 Vi 1Ω 2i, - 2 Ω + 10 V
Answer #1

Give cirait 모 ul 2A + g VA 2 노 lov cinalysis Nodal for finding V, & V₂ Using node V-10 VVy +2=0 4 V +1 + V - V2 + C (0. 9 n >ext-net in-entext 2 4 V + V₂ = 8 て 2 solving 0 & 64 svart vf=> 7= V+N2=8 7V, = 10 Y = lo 7 = 1.43V & Substinct 3 in V +V₂=8